# Coupled Sine-Gordon and $\phi^{4}$ System

### v1

@@ 3 Investigating the solutions
In Figure 4, the components of the energy density, the square of the first derivative of two functions $\phi(x)$ and $\psi(x)$ and two parts of the potential energy are plotted, to clarify the origin of the twin peaks of the energy density. Although $\phi=0$ is not a system vacuum, it has locally acted like a vacuum (note the flatness of $\phi$ around $x=0$ in Figure 3 left). According to Figure 4, the contribution related to the $\phi^{4}$ part in the energy density, ie $D_{\phi}+V_{\phi}$, is vanished around the origin.
Mass (rest energy) of the solution $MK$ is obtained as 4.1994, which is less than the average mass of $\phi^{4}$ ($0.94282\sqrt{2}/3=$ 0.9428), and Sine-Gordon kinks (8).

### v2

@@ 3 Investigating the solutions
In Figure 4, the components of the energy density, the square of the first derivative of two functions $\phi(x)$ and $\psi(x)$ and two parts of the potential energy are plotted, to clarify the origin of the twin peaks of the energy density. Although $\phi=0$ is not a system vacuum, it locally acts like a vacuum (note the flatness of $\phi$ around $x=0$ in Figure 3 left). Mass (rest energy) of the solution $MK$ is obtained as 4.1994, which is less than the average mass of $\phi^{4}$ ($0.94282\sqrt{2}/3=$ 0.9428), and Sine-Gordon kink (8).
@@ 5 Conclusions
In addition to the system vacuums, $\phi=0$ and ϕ=italic-$\phi=$ “constant” also satisfy the equations of motion (5) and (6), and is therefore a solution of the system. According to the potential (3) and Figure 1, this solution indicates a local vacuum of the system, and the potential at this solution is not an absolute minimum, but a local maximum.
The unique similarity between $D_{\psi}$ with $V_{\phi}$, and $D_{\phi}$ with $V_{\psi}$ in Figure (4) can be an inspiration for examining and analyzing the system’s solution.